OpenStudy (anonymous):

what is the cofactor expansion of 2*2 matrix

OpenStudy (smarty):

@abstracted

OpenStudy (anonymous):

So the cofactor expansion of any matrix is a method for determining the determinate of the matrix. If you have the matrix \[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\] The determinate is\[ad-bc\]

OpenStudy (anonymous):

For determinates larger than 3x3, by use of cofactor expansion, look at this http://comp.uark.edu/~jjrencis/femur/Learning-Modules/Linear-Algebra/solving/determinant/cofactor_expansion.html

OpenStudy (anonymous):

thanks ....atleast u had tried ... but i was looking for a cofactor expansion not a determinant.....

OpenStudy (anonymous):

A cofactor expansion is the exact same as the determinate. If you have the matrix \[\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\] Both the cofactor expansion and the determinate is: \[ad-bc\]

OpenStudy (anonymous):

i got it from that link u share .... thanks..@abstracted